# Landau.bib

### v1

@@ Abstract
This article is the first in a series of works on the fuzzy Landau equation, where particles interact through delocalised Coulomb collisions. Here, we establish a variational characterisation that recasts the fuzzy Landau equation within the framework of GENERIC systems (General Equations for Non-Equilibrium Reversible-Irreversible Coupling).
@@ 1 Introduction
In this paper, we consider a three-dimensional inhomogeneous (kinetic) fuzzy Landau equation:
$\partial_{t}f+v\cdot\nabla_{x}f=Q_{\sf fuz}(f,f)\quad\text{on}\quad[0,T]\times\Omega\times\mathbb{R}^{3},~{}\text{and}~{}f(0,x,v)=f_{0}(x,v).$ (1.1)
The spatial domain $\Omega$ can be $\mathbb{T}^{3}$ or $\mathbb{R}^{3}$ (we also consider this equation on $\mathbb{R}^{2d}$ for $d\geq 2$ in Section 4). This equation characterises the time evolution of the unknown $f_{t}(x,v):[0,T]\times\Omega\times\mathbb{R}^{3}\to\mathbb{R}_{+}$, describing the distribution of particles in a plasma at time $t$ at position $x$ and with velocity $v$. The initial datum $f_{0}$ is a given distribution on the phase space. The fuzzy Landau collision term is given by
$Q_{\sf fuz}(f,f)=$ $\nabla_{v}\cdot\Big{(}\int_{\Omega\times\mathbb{R}^{3}}\kappa(x-x_{*})a(v-v_{*})\big{(}f_{*}\nabla_{v}f-f\nabla_{v_{*}}f_{*}\big{)}{\,\rm d}x_{*}{\,\rm d}v_{*}\Big{)},$
where $f_{*}=f(x_{*},v_{*})$ denotes the density of particle at different position and velocity $(x_{*},v_{*})$. The kernel $a:\mathbb{R}^{3}\to\mathbb{R}^{3}$ is given by
@@
where $A:\mathbb{R}^{3}\to\mathbb{R}_{+}$ is a weight, and $\Pi_{z^{\perp}}$ denotes the projection operator onto $z^{\perp}$,
@@
In (1.3), $\gamma$ is a physical parameter, where $\gamma>0$ corresponds to the so-called hard potentials and $\gamma<0$ to the soft potentials (with a further classification of $-2\leq\gamma<0$ as the moderately soft potentials and $-4\leq\gamma<-2$ as the very soft potentials). The $\gamma=0$ and $\gamma=-3$ are known as special cases the Maxwellian and Coulomb cases respectively. We consider
$\gamma\in(-3,1].$
One relevant choice of the spatial kernel $\kappa:\mathbb{R}^{3}\to\mathbb{R}_{+}$ is $\kappa=\kappa^{\sigma}$, for a fixed parameter $\sigma\in(0,1)$, is a renormalised kernel proportional to $\exp(-\langle x/{\sqrt{\sigma}}\rangle)$, where $\langle z\rangle:=\sqrt{1+|z|^{2}}$ denotes the Japanese bracket. Various interaction kernels $A$ and spatial kernels $\kappa$ will be discussed. The precise scope can be found in Assumption 2.1 below.
By taking the limit $\sigma\to+\infty$, one recovers the homogeneous Landau equation see Remark 2.2 below. By taking the limit $\sigma\to 0$ instead, the spatial kernel $\kappa^{\sigma}$ converges to a Dirac measure. In this case, the fuzzy Landau equation (1.1) formally converges to the classical inhomogeneous Landau equation, which is given by
[…]
The classical Landau equation, which is an important partial differential equation in kinetic theory, describes the evolution of a particle distribution in a plasma, accounting for both free particle transportation and Coulomb collisions. The classical Landau collisions in (1.4) are localised in the sense that two particles with velocities $v$ and $v_{*}\in\mathbb{R}^{3}$ collide at the same position $x\in\Omega$.
In the case of the fuzzy Landau equation (1.1), we allow delocalised collisions. This means that two particles can collide with different $(x,v)$ and $(x_{*},v_{*})$, where the positions are restricted by the spatial kernel $\kappa$.
In this article, which is the first part of our series, we show that the fuzzy Landau equation (1.1) can be rigorously cast into the so-called GENERIC framework via a variational characterisation.
@@ 1.1 GENERIC structure
The GENERIC (General Equations for Non-Equilibrium Reversible Irreversible Coupling) framework provides a systematic method to derive thermodynamically consistent evolution equations describing complex systems consisting of both Hamiltonian (reversible) dynamics and gradient (irreversible) flows. It was first introduced in the context of complex fluids [GO97a, GO97b]. Over the last decade, it has proved to be a versatile modelling tool for a variety of complex systems in physics and engineering, ranging from anisotropic inelastic solids to viscoplastic solids and thermoelastic dissipative materials, to name just a few: We refer the reader to the monographs [Ott05, pavelka2018multiscale] for a detailed exposition of this framework. However, the mathematically rigorous foundations of GENERIC are still under-developed. The present work contributes towards this overarching goal.
@@
    for all functions $\mathsf{G}_{i}:\mathsf{Z}\to\mathbb{R}$, $i=1,2,3$, where the Poisson bracket {,}$\{\cdot,\cdot\}_{\mathsf{L}}$ is defined as follows:
@@
Thus, the first and second laws of thermodynamics are automatically satisfied for GENERIC systems. This, together with the structures of energy and entropy functionals strongly suggest a connection between the GENERIC framework and variational principles. However, while this idea was rigorously established for pure Hamiltonian systems (see for instance, [arnol2013mathematical]) or pure gradient-flow systems, see for instance [AGS08], it is still underdeveloped for full GENERIC systems. This is one of the key challenges in the field, see [Ottinger2018].
@@
When $\mathsf{E}=0$, the system (1.5) reduces to a gradient flow and the variational characterisation (1.7) corresponds to De Giorgi’s characterisation of gradient flows as curves of maximal slope, see for example [AGS08]. In other words, (1.7) can be viewed as an extension of De Giorgi’s formulation to GENERIC systems. A natural question is whether this can be made rigorous. This was achieved for some simple systems: The kinetic Fokker-Planck equation [DPZ13], a damped harmonic oscillator [jungel2019two, jungel2021minimizing] (these articles use a discretised version of the characterisation (1.7)), and more recently, a fuzzy Boltzmann equation [EH24]. In the following subsection, we discuss previous work on the Landau equation.
@@ 1.2 Related work
However, rigorously proving the GENERIC structure of the classical Landau equation (1.4) is a challenging problem. This is because the quadratic form in position $x$, the product $f(x,v)f(x,v_{*})$, in the classical Landau collision operator $Q_{\sf class}(f,f)$ causes the classical collision terms to not naturally be in $L^{1}(\Omega\times\mathbb{R}^{3})$, which is the natural space for the density function. In regards to solvability, within the $L^{1}$-framework, a so called renormalised solution has been studied in [Vil96], among others.
@@
$\|f(x,v)f(x_{*},v_{*})\|_{L^{1}((\Omega\times\mathbb{R}^{3})^{2})}\leq\|f(x,v)\|_{L^{1}(\Omega\times\mathbb{R}^{3})}^{2}.$ (1.8)
@@
$\partial_{t}f(v)=Q_{\sf homo}(f,f)(v),\quad f_{t}(v):\mathbb{R}^{3}\to\mathbb{R}_{+}$ (1.9)
with the homogeneous collision term $Q_{\sf homo}$ given in (2.7) below. Recently, Carrillo, Delgadino, Desvillettes and Wu [carrillo2024landau] showed that (1.9) can be interpreted as a gradient flow of the Boltzmann entropy $\mathcal{H}(f)=\int_{\mathbb{R}^{3}}f\log f{\,\rm d}v$ with respect to a certain Wasserstein-type metric.
The weak $L^{1}$ solutions of the homogeneous Landau equation (1.9) have been studied intensively for hard, Maxwellian and soft potentials [Vil98, DV00a, DV00b, Wu14, ALL15]. As a consequence of the a priori estimate (1.8), the fuzzy Landau equation (1.1) is structurally similar to the homogeneous Landau equation (1.9). In Part II of this series of articles, we will build on existing results for the homogeneous Landau equation to study the weak $L^{1}$ solutions for the fuzzy Landau equation (1.1).
The Landau equation can also be derived from the grazing limit of the Boltzmann equation, where collisions with small angular deviations play a significant role [Vil98b, He14]. The gradient flow structure of spatially homogeneous Boltzmann equation was studied in [erbar2023gradient] and a GENERIC structure for delocalised fuzzy Boltzmann equation was shown by Erbar and the second author in [EH24]. Concerning the grazing limit of the Boltzmann equation, for the spatially homogeneous case, a gradient flow perspective was discussed in [carrillo2022boltzmann]. Motivated by these results, we will in future work, study the grazing limit from the fuzzy Boltzmann equation to the fuzzy Landau equation using their respective GENERIC variational characterisations in [EH24] and the results of this present work.
Delocalised collision models have previously been studied in other contexts, such as the Enskog equation for high-density large diameter particle systems [Ark86, AC89], more general delocalised kinetic models recently discussed in [FZF24], and an optimal transport perspective was discussed in [GJF25]. Finally, related discrete variational formulations, akin to the celebrated Jordan–Kinderlehrer–Otto [JKO98] scheme, for non-gradient flow systems were studied in [gigli2012variational, duong2014conservative, duong2018fundamental, adams2022operator, duong2022entropic], and extensions of symplectic integrators for Hamiltonian systems to GENERIC systems were discussed in [ottinger2018generic, shang2020structure].
@@ 1.3 Variational characterisation
mass $\int_{\Omega\times\mathbb{R}^{3}}f_{t}{\,\rm d}x{\,\rm d}v=\int_{\Omega\times\mathbb{R}^{3}}f_{0}{\,\rm d}x{\,\rm d}v,$
momentum $\int_{\Omega\times\mathbb{R}^{3}}f_{t}v{\,\rm d}x{\,\rm d}v=\int_{\Omega\times\mathbb{R}^{3}}f_{0}v{\,\rm d}x{\,\rm d}v,$
energy $\int_{\Omega\times\mathbb{R}^{3}}f_{t}|v|^{2}{\,\rm d}x{\,\rm d}v=\int_{\Omega\times\mathbb{R}^{3}}f_{0}|v|^{2}{\,\rm d}x{\,\rm d}v.$
We define the Boltzmann entropy functional as $\mathcal{H}(f)=\int_{\Omega\times\mathbb{R}^{3}}f\log f{\,\rm d}x{\,\rm d}v$. Formally, the entropy identity holds
@@
$\mathcal{D}(f)=\frac{1}{2}\int_{(\Omega\times\mathbb{R}^{3})^{2}}\kappa ff_{*}\big{|}\tilde{\nabla}\log f\big{|}^{2}{\,\rm d}\eta\geq 0,$ (1.10)
where throughout the paper we use the shortened notation ${\,\rm d}\eta={\,\rm d}x_{*}{\,\rm d}v_{*}{\,\rm d}x{\,\rm d}v$ to denote the Lebesgue measure on $\mathbb{R}^{12}$ and $\tilde{\nabla}$ is the fuzzy Landau gradient operator defined below.
@@
$\mathsf{E}(f)=\int_{\Omega\times\mathbb{R}^{3}}\frac{|v|^{2}}{2}f{\,\rm d}x{\,\rm d}v\quad\text{and}\quad\mathsf{S}(f)=-\mathcal{H}(f),$
@@
$\mathsf{M}(f)g=-\frac{1}{2}\tilde{\nabla}\cdot\Big{(}\kappa ff_{*}\tilde{\nabla}g\Big{)}\quad\text{and}$
@@
where we follow [carrillo2024landau], to define the fuzzy Landau gradient
~$\tilde{\nabla}f(x,v)=\sqrt{A}\Pi_{(v-v_{*})^{\perp}}\big{(}\nabla_{v}f(x,v)-\nabla_{v_{*}}f(x_{*},v_{*})\big{)}.$ (1.11)
Note that $A=A(v-v_{*})$ and for the homogeneous Landau equation, $A=|v-v_{*}|^{2+\gamma}$. The corresponding divergence, $\tilde{\nabla}\cdot$, is defined via integration by parts for a function $B(x,x_{*},v,v_{*}):(\Omega\times\mathbb{R}^{3})^{2}\to\mathbb{R}^{3}$
$\int_{(\Omega\times\mathbb{R}^{3})^{2}}\tilde{\nabla}g\cdot B{\,\rm d}\eta=-\int_{\Omega\times\mathbb{R}^{3}}g\tilde{\nabla}\cdot B{\,\rm d}v{\,\rm d}x\quad\forall\,g:\Omega\times\mathbb{R}^{3}\to\mathbb{R}.$ (1.12)
@@
~$\tilde{\nabla}\cdot B=\nabla_{v}\cdot\Big{(}\int_{\Omega\times\mathbb{R}^{3}}\sqrt{A}\Pi_{(v-v_{*})^{\perp}}\big{(}B(x,x_{*},v,v_{*})-B(x_{*},x,v_{*},v)\big{)}{\,\rm d}x_{*}{\,\rm d}v_{*}\Big{)}.$
@@
$\partial_{t}f+v\cdot\nabla_{x}f+\frac{1}{2}\tilde{\nabla}\cdot U_{t}=0,$ (1.13)
where $U:[0,T]\times(\Omega\times\mathbb{R}^{3})^{2}\to\mathbb{R}$ denotes the grazing rate (following the same terminology as in the homogeneous case [carrillo2024landau]). Notice that, when $U_{t}=-\kappa ff_{*}\tilde{\nabla}\log f$ equation (1.13) recovers the fuzzy Landau equation (1.1). We also define the following curve action associated to (1.13) (see in Definition 3.2 and Lemma 3.3 in Section 3.2.2 for the precise definition and characterisation):
$\mathcal{A}(f,U)=\frac{1}{2}\int_{(\Omega\times\mathbb{R}^{3})^{2}}\frac{|U|^{2}}{ff_{*}\kappa}{\,\rm d}\eta.$
@@
Let the kernels $(A,\kappa)$ satisfy Assumption 2.1. Let $(f_{t},\mathcal{U}_{t})$ be solutions to the transport grazing rate equation (1.13). Let $(f_{t})_{t\in[0,T]}$ be a curve of probability density on $\Omega\times\mathbb{R}^{3}$ such that $\int_{\Omega\times\mathbb{R}^{3}}(\langle x\rangle^{2}+\langle v\rangle^{2+|\gamma|})f_{t}{\,\rm d}x{\,\rm d}v<+\infty$ for any $t\in[0,T]$, initial entropy $|\mathcal{H}(f_{0})|<+\infty$, and satisfies Assumption 2.2 below.
@@
[…]
Fuzzy Landau equation on $\mathbb{R}^{2d}$ for $d\geq 2$. We note that while [carrillo2024landau] is restricted to three dimensions due to the use of a three-dimensional cross-product identity (see Section 2.4), the variational characterisation in Theorem 1.1 remains valid in all dimensions ($d\geq 2$) for the general class of kernels $(\bar{A},\kappa)$, which has been considered for the spatially homogeneous Boltzmann equation and the kinetic fuzzy Boltzmann equation [erbar2023gradient, EH24]. For precise statements, see Remark 2.1 and Section 4 (in particular Theorem 4.1). Our results also open up a couple of interesting avenues for future research, which will be discussed further in Section 5.
The rest of the paper is organised as follows. In Section 2, we introduce the necessary notations and discuss curves that satisfy Assumption 2.2. In Section 3, we perform regularisations and error estimates to rigorously verify the chain rue and prove our main result, Theorem 1.1. In Section 4, we discuss the fuzzy Landau equations with a general class of kernels that the main Theorem 1.1 holds for all dimensions $(d\geq 2)$. Further discussions on future research are presented in Section 5. Finally, some detailed and technical computations are given in Appendices.

### v2

# On a fuzzy Landau Equation: Part I. A variational approach

@@ Abstract
In this paper, we introduce a fuzzy Landau equation, where particles interact through delocalised Coulomb collisions. We establish a variational characterisation that recasts the fuzzy Landau equation within the framework of GENERIC systems (General Equations for Non-Equilibrium Reversible-Irreversible Coupling), obtaining an equivalent characterisation of $\mathcal{H}$-solutions.
@@ 1 Introduction
In this paper, we consider the inhomogeneous (kinetic) fuzzy Landau equation:
$\partial_{t}f+v\cdot\nabla_{x}f=Q_{\sf fuz}(f,f)\quad\text{on}\quad[0,T]\times\Omega\times\mathbb{R}^{d},~\text{and}~f(0,x,v)=f_{0}(x,v).$ (1.1)
The spatial domain $\Omega$ can be $\mathbb{T}^{d}$ or $\mathbb{R}^{d}$ ($d\geq 2$). This equation characterises the time evolution of the unknown $f_{t}(x,v):[0,T]\times\Omega\times\mathbb{R}^{d}\to\mathbb{R}_{+}$, describing the distribution of particles in a plasma at time $t$ at position $x$ and with velocity $v$. The initial datum $f_{0}$ is a given distribution on the phase space. The fuzzy Landau collision term is given by
$Q_{\sf fuz}(f,f)=$ $\nabla_{v}\cdot\Big(\int_{\Omega\times\mathbb{R}^{d}}\kappa(x-x_{*})a(v-v_{*})\big(f_{*}\nabla_{v}f-f\nabla_{v_{*}}f_{*}\big){\,\rm d}x_{*}{\,\rm d}v_{*}\Big),$
where $f_{*}=f(x_{*},v_{*})$ denotes the density of particle at different position and velocity $(x_{*},v_{*})$. The kernel $a:\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ is given by
@@
where $A:\mathbb{R}^{d}\to\mathbb{R}_{+}$ is a weight, and $\Pi_{z^{\perp}}$ denotes the projection operator onto $z^{\perp}$,
@@
In (1.3), $\gamma$ is a physical parameter, where $\gamma>0$ corresponds to the so-called hard potentials and $\gamma<0$ to the soft potentials (with a further classification of $2<\gamma<0$ as the moderately soft potentials and $d<\gamma\leq-2$ as the very soft potentials). In particular, the $\gamma=0$ is known as the Maxwellian case while the case $\gamma=-d$ is known as the Coulomb case. In this paper, we consider
$\gamma\in(-\gamma_{d},1],$
where $\gamma_{d}=\min(d,4)$. The condition $\gamma>-d$ ensures the local integrability of $|z|^{\gamma}$, and the solutions to the fuzzy Landau equations are well-defined for $\gamma>-4$, as shown in Part II [DH25].
One relevant choice of the spatial kernel $\kappa:\mathbb{R}^{d}\to\mathbb{R}_{+}$ is $\kappa=\kappa^{\sigma}$, where for a fixed parameter $\sigma\in(0,1)$, $\kappa^{\sigma}$ is a renormalised kernel proportional to $\exp(-\langle x/{\sqrt{\sigma}}\rangle)$, where $\langle z\rangle:=\sqrt{1+|z|^{2}}$ denotes the Japanese bracket. With this choice of the spatial kernel, the fuzzy Landau equation can be viewed as an intermediate model between the spatially homogeneous Landau equation and the classical inhomogeneous Landau equation. In fact, by taking the limit $\sigma\to+\infty$, one recovers the homogeneous Landau equation, see Remark 2.4 below. By taking the limit $\sigma\to 0$ instead, the spatial kernel $\kappa^{\sigma}$ converges to a Dirac measure. In this case, the fuzzy Landau equation (1.1) formally converges to the classical inhomogeneous Landau equation, which is given by
$\left\{\begin{aligned} &\partial_{t}f+v\cdot\nabla_{x}f=Q_{\sf class}(f,f)\\ &Q_{\sf class}(f,f)=\nabla_{v}\cdot\Big(\int_{\mathbb{R}^{d}}a\big(f(x,v_{*})\nabla_{v}f(x,v)-f(x,v)\nabla_{v_{*}}f(x,v_{*})\big){\,\rm d}v_{*}\Big).\end{aligned}\right$. (1.4)
The classical Landau equation, which is an important partial differential equation in kinetic theory, describes the evolution of a particle distribution in a plasma, accounting for both free particle transportation and Coulomb collisions. The classical Landau collisions in (1.4) are localised in the sense that two particles with velocities $v$ and $v_{*}\in\mathbb{R}^{d}$ collide at the same position $x\in\Omega$.
In the case of the fuzzy Landau equation (1.1), we allow delocalised collisions. This means that two particles can collide with different $(x,v)$ and $(x_{*},v_{*})$, where the positions are restricted by the spatial kernel $\kappa$. Various interaction kernels $A$ and spatial kernels $\kappa$ will be discussed. The precise scope can be found in Assumption 2.1 below.
In this article, which is the first part of our series, we show that the fuzzy Landau equation (1.1) can be rigorously cast into the so-called GENERIC framework via a variational formulation, obtaining an equivalent characterisation of $\mathcal{H}$-solutions.
@@ 1.1 GENERIC structure
The GENERIC (General Equations for Non-Equilibrium Reversible Irreversible Coupling) framework provides a systematic method to derive thermodynamically consistent evolution equations describing complex systems consisting of both Hamiltonian (reversible) dynamics and gradient (irreversible) flows. It was first introduced in the context of complex fluids [GÖ97, GÖ97a]. Over the last decade, it has proved to be a versatile modelling tool for a variety of complex systems in physics and engineering, ranging from anisotropic inelastic solids to viscoplastic solids and thermoelastic dissipative materials, to name just a few: We refer the reader to the monographs [Ött05, PKG18] for a detailed exposition of this framework. However, a mathematically rigorous foundation for GENERIC is still under-developed. The present work contributes towards this overarching goal.
@@
    for all functions $\mathsf{G}_{i}:\mathsf{Z}\to\mathbb{R}$, $i=1,2,3$, where the Poisson bracket $\{\cdot,\cdot\}_{\mathsf{L}}$ is defined as follows:
@@
Thus, the first and second laws of thermodynamics are automatically satisfied for GENERIC systems. This, together with the structures of energy and entropy functionals strongly suggest a connection between the GENERIC framework and variational principles. However, while this idea was rigorously established for pure Hamiltonian systems (see for instance, [Arn13]) or pure gradient-flow systems, see for instance [AGS08], it is still underdeveloped for full GENERIC systems. This is one of the key challenges in the field, see [Ött18a].
@@
When $\mathsf{E}=0$, the system (1.5) reduces to a gradient flow and the variational characterisation (1.7) corresponds to De Giorgi’s characterisation of gradient flows as curves of maximal slope, see for example [AGS08]. In other words, (1.7) can be viewed as an extension of De Giorgi’s formulation to GENERIC systems. A natural question is whether this can be made rigorous. This was achieved for some simple systems: The kinetic Fokker-Planck equation [DPZ13], a damped harmonic oscillator [JST19, JST21] (these articles use a discretised version of the characterisation (1.7)), and more recently, a fuzzy Boltzmann equation [EH25]. In the following subsection, we discuss previous work on the Landau equation.
@@ 1.2 Related work
However, rigorously proving the GENERIC structure of the classical Landau equation (1.4) is a challenging problem. This is because the quadratic form in position $x$, the product $f(x,v)f(x,v_{*})$, in the classical Landau collision operator $Q_{\sf class}(f,f)$ causes the classical collision terms to not naturally be in $L^{1}(\Omega\times\mathbb{R}^{d})$, which is the natural space for the density function. In regards to solvability, within the $L^{1}$-framework, a so called renormalised solution has been studied in [Vil96], among others.
@@
$\|f(x,v)f(x_{*},v_{*})\|_{L^{1}((\Omega\times\mathbb{R}^{d})^{2})}\leq\|f(x,v)\|_{L^{1}(\Omega\times\mathbb{R}^{d})}^{2}.$ (1.8)
@@
$\partial_{t}f(v)=Q_{\sf homo}(f,f)(v),\quad f_{t}(v):\mathbb{R}^{d}\to\mathbb{R}_{+}$ (1.9)
with the homogeneous collision term $Q_{\sf homo}$ given in (2.8) below. Recently, Carrillo, Delgadino, Desvillettes and Wu [CDDW24] showed that (1.9) can be interpreted as a gradient flow of the Boltzmann entropy $\mathcal{H}(f)=\int_{\mathbb{R}^{d}}f\log f{\,\rm d}v$ with respect to a certain Wasserstein-type metric.
The weak $L^{1}$ solutions of the homogeneous Landau equation (1.9) have been studied intensively for hard, Maxwellian and soft potentials [Vil98, DV00, DV00a, Wu14, ALL15]. As a consequence of the a priori estimate (1.8), the fuzzy Landau equation (1.1) is structurally similar to the homogeneous Landau equation (1.9). In Part II of this series of articles, we will build on existing results for the homogeneous Landau equation to study the weak $L^{1}$ solutions for the fuzzy Landau equation (1.1).
The Landau equation can also be derived from the grazing limit of the Boltzmann equation, where collisions with small angular deviations play a significant role [Vil98a, He14]. The gradient flow structure of spatially homogeneous Boltzmann equation was studied in [Erb23] and a GENERIC structure for delocalised fuzzy Boltzmann equation was shown by Erbar and the second author in [EH25]. Concerning the grazing limit of the Boltzmann equation, for the spatially homogeneous case, a gradient flow perspective was discussed in [CDW22]. Motivated by these results, in part III [DGH25] we study the grazing limit from the fuzzy Boltzmann equation to the fuzzy Landau equation using their respective GENERIC variational characterisations in [EH25] and the results of this present work.
Delocalised collision models have previously been studied in other contexts, such as the Enskog equation for high-density large diameter particle systems [Ark86, AC89], more general delocalised kinetic models recently discussed in [CCG24], and an optimal transport perspective was discussed in [BMQ25]. Finally, related discrete variational formulations, akin to the celebrated Jordan–Kinderlehrer–Otto [JKO98] scheme, for non-gradient flow systems were studied in [GM12, DPZ14, DT18, ADR22, DAR22], and extensions of symplectic integrators for Hamiltonian systems to GENERIC systems were discussed in [Ött18, SÖ20].
@@ 1.3 Variational characterisation
mass $\int_{\Omega\times\mathbb{R}^{d}}f_{t}{\,\rm d}x{\,\rm d}v=\int_{\Omega\times\mathbb{R}^{d}}f_{0}{\,\rm d}x{\,\rm d}v,$
momentum $\int_{\Omega\times\mathbb{R}^{d}}f_{t}v{\,\rm d}x{\,\rm d}v=\int_{\Omega\times\mathbb{R}^{d}}f_{0}v{\,\rm d}x{\,\rm d}v,$
energy $\int_{\Omega\times\mathbb{R}^{d}}f_{t}|v|^{2}{\,\rm d}x{\,\rm d}v=\int_{\Omega\times\mathbb{R}^{d}}f_{0}|v|^{2}{\,\rm d}x{\,\rm d}v.$
We define the Boltzmann entropy functional as $\mathcal{H}(f)=\int_{\Omega\times\mathbb{R}^{d}}f\log f{\,\rm d}x{\,\rm d}v$. Formally, the entropy identity holds
@@
$\mathcal{D}(f)=\frac{1}{2}\int_{(\Omega\times\mathbb{R}^{d})^{2}}\kappa ff_{*}\big|\widetilde{\nabla}\log f\big|^{2}{\,\rm d}\eta\geq 0,$ (1.10)
where throughout the paper we use the shortened notation ${\,\rm d}\eta={\,\rm d}x_{*}{\,\rm d}v_{*}{\,\rm d}x{\,\rm d}v$ to denote the Lebesgue measure on $\mathbb{R}^{4d}$ and $\widetilde{\nabla}$ is the fuzzy Landau gradient operator defined below.
@@
$\mathsf{E}(f)=\int_{\Omega\times\mathbb{R}^{d}}\frac{|v|^{2}}{2}f{\,\rm d}x{\,\rm d}v\quad\text{and}\quad\mathsf{S}(f)=-\mathcal{H}(f),$
@@
$\mathsf{M}(f)g=-\frac{1}{2}\widetilde{\nabla}\cdot\Big(\kappa ff_{*}\widetilde{\nabla}g\Big)\quad\text{and}$
@@
where we follow [CDDW24], to define the fuzzy Landau gradient
$\widetilde{\nabla}f(x,v)=\sqrt{A}\Pi_{(v-v_{*})^{\perp}}\big(\nabla_{v}f(x,v)-\nabla_{v_{*}}f(x_{*},v_{*})\big).$ (1.11)
Note that $A=A(v-v_{*})$ and for the homogeneous Landau equation, $A=|v-v_{*}|^{2+\gamma}$. The corresponding divergence, $\widetilde{\nabla}\cdot$, is defined via integration by parts for a function $B(x,x_{*},v,v_{*}):(\Omega\times\mathbb{R}^{d})^{2}\to\mathbb{R}^{d}$
$\int_{(\Omega\times\mathbb{R}^{d})^{2}}\widetilde{\nabla}g\cdot B{\,\rm d}\eta=-\int_{\Omega\times\mathbb{R}^{d}}g\widetilde{\nabla}\cdot B{\,\rm d}v{\,\rm d}x\quad\forall\,g:\Omega\times\mathbb{R}^{d}\to\mathbb{R}.$ (1.12)
@@
$\widetilde{\nabla}\cdot B=\nabla_{v}\cdot\Big(\int_{\Omega\times\mathbb{R}^{d}}\sqrt{A}\Pi_{(v-v_{*})^{\perp}}\big(B(x,x_{*},v,v_{*})-B(x_{*},x,v_{*},v)\big){\,\rm d}x_{*}{\,\rm d}v_{*}\Big).$
@@
$\partial_{t}f+v\cdot\nabla_{x}f+\frac{1}{2}\widetilde{\nabla}\cdot U_{t}=0,$ (1.13)
where $U:[0,T]\times(\Omega\times\mathbb{R}^{d})^{2}\to\mathbb{R}$ denotes the grazing rate (following the same terminology as in the homogeneous case [CDDW24]). Notice that, when $U_{t}=-\kappa ff_{*}\widetilde{\nabla}\log f$ equation (1.13) recovers the fuzzy Landau equation (1.1). We also define the following curve action associated to (1.13) (see in Definition 3.2 and Lemma 3.3 in Section 3.2.2 for the precise definition and characterisation):
$\mathcal{A}(f,U)=\frac{1}{2}\int_{(\Omega\times\mathbb{R}^{d})^{2}}\frac{|U|^{2}}{ff_{*}\kappa}{\,\rm d}\eta.$
@@
Let the kernels $(A,\kappa)$ satisfy Assumption 2.1. Let $(f_{t},\mathcal{U}_{t})$ be solutions to the transport grazing rate equation (1.13). Let $(f_{t})_{t\in[0,T]}$ be a curve of probability density on $\Omega\times\mathbb{R}^{d}$ such that $\int_{\Omega\times\mathbb{R}^{d}}(\langle x\rangle^{2}+\langle v\rangle^{2+|\gamma|})f_{t}{\,\rm d}x{\,\rm d}v<+\infty$ for any $t\in[0,T]$, initial entropy $|\mathcal{H}(f_{0})|<+\infty$, and satisfies Assumption 2.7 below.
@@
The definition of $\mathcal{H}$-solutions is provided in Definition 2.5, following a parallel approach to the $\mathcal{H}$-solutions framework for homogeneous Landau equation introduced by [Vil98a]. The proof of Theorem 1.1, which will be given in Section 3 (it is stated as Theorem 3.16 there), amounts to rigorously verifying a chain rule for the transport grazing rate equation (1.13). We achieve this using regularisation techniques and error estimates. Our analysis requires significant technical improvements compared to the homogeneous case in [CDDW24] to account for the spatial variable $x$ and the transport term $v\cdot\nabla_{x}f(x,v)$. We note that while [CDDW24] is restricted to three dimensions, the variational characterisation in Theorem 1.1 remains valid in all dimensions ($d\geq 2$). Our results also open up a couple of interesting avenues for future research, which will be discussed further in Section 4.
The rest of the paper is organised as follows. In Section 2, we introduce the necessary notations and discuss curves that satisfy Assumption 2.7. In Section 3, we perform regularisations and error estimates to rigorously verify the chain rue and prove our main result, Theorem 1.1. In Section 4, we discuss various GENERIC building blocks for fuzzy Landau equations and possible directions for future works. Finally, some detailed and technical computations are given in Appendices.
This article is the first in a series of works on the fuzzy Landau equation. The second paper [DH25] is devoted to show the existence and regularity results of fuzzy Landau equations. The third paper [DGH25] studies the grazing collision limit of fuzzy Boltzmann equations to fuzzy Landau equations by using the variational method introduced in this paper.